answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the meaning of the word likelihood?

In statistics, a likelihood function (often simply likelihood) is a function of a statistical model. The likelihood of a set parameter values, given outcomes x, is equal to the probability of those observed outcome.

What is the probability of landing on 4 on a spinner?

depends on how many 4's and how many total spaces, assuming they are all equal size. If not then it depends on area of each sector.

When two dice are rolled what is the probability of getting a sum greater than or equal to 9?

There are 36 different outcomes when rolling two dice.

4 combinations make 9.

3 make 10, 2 make 11 and 1 makes 12.

4+3+2+1 = 10

10 out of 36 = 10/36 = 0.277777 = 27.8 %

What is the probability of picking a card that will be a diamond or the ace of spades?

If you mean in one draw from a standard 52- card deck, there are 13 diamonds and ace of spades have 14 chances out of 52 for a probability of 14/52 or about 26.6923 %

There are nine pong balls numbered 1-9 in a bag If three balls are randomly selected without replacement what is the probability expressed as a common fraction that the sum of the number be odd?

10/21

First, observe that for the sum of the three balls to be odd, either one is odd and two are even, or all three are odd. Since the sum of two odd numbers is even, the sum of two even number is even, and the sum of an even number and an odd number is odd.

P(exactly one is odd)=(5/9)*(4/8)*(3/7)*3=180/504=5/14

P(all three are odd)=(5/9)*(4/8)*(3/7)=60/504=5/42

P(sum is odd)=5/14+5/42=15/42+5/42=20/42=10/21, which is approximately .48

What is the probability of winning the jackpot on the slot machines?

It depends on the machine, but hitting the multi-million dollar progressive jackpot on the Megabucks machine is almost 50 million to one. In comparison, picking 6 out of 49 numbers in the lottery is about 14 million to one.

Can a union steward be a supervisor?

No, Stewards represents employees with work place issues. The supervisior hires. fires. disaplines and evalutes. Issues in the work place are because of actions of a supervisor.

Why you add 273 to convert celsius to kelvin?

The answer is 546.15 K (approx.). The Kelvin scale starts at absolute zero and is used in scientific laboratories. Celsius is for general use and set 0 and 100 as melting and boiling point of water respectively. To convert from C to K, add 273.15 to C.

How you get the mode by applying probability distribution?

The mode is the most probable value. Often, you determine the mode by plotting the experimental probability distribution, and finding the peak value. The mode is not necessarily the same as the mean nor the median, unless the distribution is symmetrical.

How to calcalute the probability of 1 through 12?

There is insufficient information in the question to answer it. Please restate the question.

How many ways can you choose even cards from a deck of 52?

2,4,6,8,10 are even cards of a deck. There are 52 cards in all. Even cards may be chosen in 52C5 = 52x51x50x49x48 / 5x4x3x2x1 = 2,598,960 ways.

What are more rare fraternal or identical twins?

i say that identical twins are the least rare because like me a fraternal twin i only know 1 pair of fraternal twin.






A card is drawn from a standard deck of 52 playing cards Find the probability that the card is an ace or a black card?

This is a probability case of statistics using the addition rule

P(A or B) = P(A) + P(B) - P(A and B).

The output can happen at the same time, for that reason it is not mutually exclusive. The output can be an ace and a black card.

first case: how many ace's we got=4 over the number of total cards=52

second case: how many black cards=26 over the total number of cards=52

BOTH cases: they can be just 2 cards that can be aces AND black.

(4/52)+(26/52)-(2/52)= 7/13=.538461538462= .5385

How do you do probabilities?

Example of probability is if you have 10 blue gumballs 5 pink gumballs and 3 orange gumballs the answer is 10 because the more gumballs you have is the one you will pick without looking

if you still dont understand please comment below,thanks for all questions:)

When a single card is drawn from an ordinary 52 card deck what is the probability that you draw a black queen?

As there are 2 black queens in a deck, the chances of pulling one from a 52 card deck is 2/52 (or 1/26).

What is the probability of rolling a 4 and a 1 on a 6 sided die?

The answer depends on whether or not order is important:

If you want a 4 first, then a 1, the probability is (1/6) x (1/6) = (1/36) =~ 0.028.

If you don't care whether the four or the one comes first, then the probability is 2 * (1/36) =~ 0.056.

You are having a baby how do you know if its a boy or girl?

go to the doctor and get an ultrasound, then the doctor can tell you if its a boy or a girl

What is a population in statistics?

The population is the set of all units which are being studied. They need not be people: for example if I am trying to find the most popular colour for cars, the population would be all cars. If I was looking at the density of earthworms in my garden, my population could be each 20 cm x 20 cm in my garden.